There are two white, one blue and one red balloon.
Petya has guessed a colour.
Vasya tries to guess the given colour at random. What is the probability that Vasya will guess?
There is one colour out of three. It doesn't matter how many balls there are - the colour is chosen. If Vasya picks the first one out of the three, the probability of guessing it is 1/3. But I don't understand, what's the trick of the game? Is it that the number of balls is mixed in with the number of colours, to confuse them?
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and + 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the point of the problem? It's an elementary problem.
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the point of the problem? It's an elementary problem.
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and + 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the joke of the problem? It's an elementary problem.
Allow the use of cookies to log in to the MQL5.com website.
Please enable the necessary setting in your browser, otherwise you will not be able to log in.
You are signed in. Two-factor authentication has been disabled. Set it up again and scan a new QR code, or continue with two-factor authentication disabled.
There are two white, one blue and one red balloon.
Petya has guessed a colour.
Vasya tries to guess the given colour at random. What is the probability that Vasya will guess?
There is one colour out of three. It doesn't matter how many balls there are - the colour is chosen. If Vasya picks the first one out of the three, the probability of guessing it is 1/3. But I don't understand, what's the trick of the game? Is it that the number of balls is mixed in with the number of colours, to confuse them?
There are two white, one blue and one red balloons.
Petya has puzzled a colour.
Vasya tries to guess the given colour at random. What is the probability that Vasya will guess?
It depends on what he gets for it. Quite seriously, by the way ...
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and + 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the point of the problem? It's an elementary problem.
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the point of the problem? It's an elementary problem.
The situation will be different if Vasi has the balls in an opaque bag and that bag serves as a random colour generator. Then the probability changes. 2 white + 1 blue and + 1 red = 4 balls. The probability of white falling out = 2/4 = 1/2 = 50%. Probability of guessing blue = probability of guessing red = 1/4 = 25 per cent.
So what's the joke of the problem? It's an elementary problem.
So they're in the opaque one :)
X^X^ ... ^X =2.7182818285
Solve the equation.
Top down. Brackets from top to bottom